# Indecomposable involutive set-theoretical solutions to the Yang–Baxter equation of size $p^{2}$

### v2

@@ 5. Summary
    The parameters $\alpha,\Phi$ and $\alpha^{\prime},\Phi^{\prime}$ define isomorphic cycle sets if and only if $\alpha=\alpha^{\prime}$ and there is a $\beta\in\mathbb{Z}_{p}^{\ast}$ such that $\beta^{-1}\Phi(\beta x)=\Phi^{\prime}(x)$ for all $x\in\mathbb{Z}_{p}$.
@@
2.4 tells us that the indecomposable cycle sets of size $p^{2}$ that have finite multipermutation level $1$ and $2$, are exactly the ones described in (1) resp. (2). On the other hand, the irretractable cycle sets are classified, up to isomorphism in 4.8 and make up case (3). ∎
5.1. Indecomposable set-theoretic solutions of size $p^{2}$
@@
In fact, these solutions are isomorphic to those constructed in [6, Theorem 5.1] as we will show in the remainder of this section. In particular, this answers [6, Question 7.4] affirmatively.
@@
Similarly, we will denote the solution constructed in 5.2 with parameters $f,t$ as $r^{f,t}$ with first component
@@
It remains to show that, with the conditions for $\Phi$ and $\alpha$ given in 5.1, the parameters $f=f_{\Phi,\alpha}$ and $t=\alpha^{-1}$ satisfy the properties required by 5.2. Since $\Phi$ satisfies (S1) and (S3), so does $f_{\Phi,\alpha}$. Moreover, since $\Phi(\alpha i)=\alpha\Phi(i)$, we have that $f_{\Phi,\alpha}(\alpha^{-s}i)=-\Phi(\alpha\alpha^{-s}i)=-\alpha^{-s}\Phi(\alpha i)=\alpha^{-s}f_{\Phi,\alpha}(i)$. Hence $f_{\Phi,\alpha}$ satisfies (S2) if and only if $(\alpha^{-s}-1)f_{\Phi,\alpha}(0)=0$ for all $s\in\mathbb{Z}$, which is equivalent to $(\alpha-1)\Phi(0)=0$. But the latter is a consequence of the properties of $\Phi$ and $\alpha$ as $\Phi(0)=\Phi(\alpha 0)=\alpha\Phi(0)$.
@@ Acknowledgements
The authors would also like to express their grattitude to Marco Castelli for pointing out that one results in an earlier version of the manuscript was already proved by Rump.

### v3

@@ 5. Summary
    The parameters $\alpha,\Phi$ and $\alpha^{\prime},\Phi^{\prime}$ define isomorphic cycle sets if and only if $\alpha=\alpha^{\prime}$ and there is a $\beta\in\mathbb{Z}_{p}^{\ast}$ such that $\beta^{-1}\Phi(\beta x)=\Phi^{\prime}(x)$, for all $x\in\mathbb{Z}_{p}$.
@@
2.4 tells us that the indecomposable cycle sets of size $p^{2}$ that have finite multipermutation level $1$ and $2$, are exactly the ones described in (1) resp. (2). On the other hand, the irretractable cycle sets are classified, up to isomorphism in 4.7 and make up case (3). ∎
Using the correspondence between cycle sets and set-theoretical solutions, the previous theorem, reformulated in terms of set-theoretical solutions, is the following.
Corollary 5.2.
Each non-degenerate, indecomposable, involutive set-theoretical solution $(X,r)$ to the Yang–Baxter equation of size $p^{2}$ for some prime $p$ is isomorphic to one of the following solutions:
  1. (1)
    $X=\mathbb{Z}_{p^{2}}$, with $r(x,y)=(y+1,x-1)$.
  2. (2)
    $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, with
    $\chi_{0}(x)$ =$=\begin{cases}1&x=0\\ 0&x\neq 0\end{cases}.$
  3. (3)
    $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, with
5.1. Indecomposable set-theoretical solutions of size $p^{2}$
@@
In fact, these solutions are isomorphic to those constructed in [6, Theorem 5.1] as we will show in the remainder of this section. In particular, this answers [6, Question 7.3] affirmatively.
@@
Similarly, we will denote the solution constructed in 5.3 with parameters $f,t$ as $r^{f,t}$ with first component
@@
It remains to show that, with the conditions for $\Phi$ and $\alpha$ given in 5.1, the parameters $f=f_{\Phi,\alpha}$ and $t=\alpha^{-1}$ satisfy the properties required by 5.3. Since $\Phi$ satisfies (S1) and (S3), so does $f_{\Phi,\alpha}$. Moreover, since $\Phi(\alpha i)=\alpha\Phi(i)$, we have that $f_{\Phi,\alpha}(\alpha^{-s}i)=-\Phi(\alpha\alpha^{-s}i)=-\alpha^{-s}\Phi(\alpha i)=\alpha^{-s}f_{\Phi,\alpha}(i)$. Hence $f_{\Phi,\alpha}$ satisfies (S2) if and only if $(\alpha^{-s}-1)f_{\Phi,\alpha}(0)=0$ for all $s\in\mathbb{Z}$, which is equivalent to $(\alpha-1)\Phi(0)=0$. But the latter is a consequence of the properties of $\Phi$ and $\alpha$ as $\Phi(0)=\Phi(\alpha 0)=\alpha\Phi(0)$.
@@ Acknowledgements
The authors would also like to express their gratitude to Marco Castelli for pointing out that one results in an earlier version of the manuscript was already proved by Rump.